I got this question and may you please answer as soon as possible.
A cylindrical water tank is 200 cm high and has radius of 70 cm.
a) Calculate the volume of the tank in, if 'Pi equals 22 over 7 (Pi = 22/7)'
b) If every 1000 cm3 (1000 cm cubed) holds 1 litre of water, how many litres will the tank hold?
Thank you!
~official_bsanielle

Answer by lwsshak3(11628)

I got this question and may you please answer as soon as possible.
A cylindrical water tank is 200 cm high and has radius of 70 cm.
a) Calculate the volume of the tank in, if 'Pi equals 22 over 7 (Pi = 22/7)'
b) If every 1000 cm3 (1000 cm cubed) holds 1 litre of water, how many litres will the tank hold?
***
volume=area of end*height
area of end=π*radius^2=π*70^2=4900 π cm^2
volume=4900π*200=980000π cm^3=980π litres=980(22/7)=3080 litres
tank will hold 3080 litres